OpenStudy (anonymous):

Find the slope of the line that contains the following points. R(-3, 5), S(3, -2)

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

Just helped. Use my former method.

OpenStudy (anonymous):

slope=y2-y1/x2-x1 ---->slope=-2-5/3--3 --->slope=-7/6